OpenStudy (anonymous):

e^x e^(x+1)=1 how do you combine the exponential expressions in a single exponential expression

OpenStudy (anonymous):

\[x^a\cdot x^b = x^{a+b}\]

OpenStudy (anonymous):

\[ a^n\times a^m = a^{n+m} \]In this case \(a=e, n=x, m=x+1\).

OpenStudy (anonymous):

so e^x*e^x=e^2x

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

actually wouldnt it be e^x*e^x=e^2x+1

OpenStudy (anonymous):

if you wrote it correctly \[e^x\cdot e^{x+1} = e^{2x+1}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so what's x?

OpenStudy (anonymous):

so first I would have to convert it into a logarithmic equation?

OpenStudy (anonymous):

what is 1?

OpenStudy (anonymous):

\[1=e^0\]